# Flatbands in tight-binding lattices with anisotropic potentials

### v1

@@ IV.2 Arctan potential
Figure 7: Subfigures (a) and (b): The band structure around the flatband at $E=0$, obtained by diagonalizing (12) and (III), respectively. The potential (19) parameters are $\lambda=4$(a) and $\lambda=3$(b), $\xi=5$, and system size $L=201$ along $z$.
We consider the following trigonometric potential in Hamiltonians (12) and (III):
$V(z)=\lambda\arctan\left(\frac{z}{\xi}\right)\;.$ (19)
@@
with $\xi=5$, for two different directions $(1,1)$ with $z=n_{1}+n_{2}$ and $(3,1)$ with $z=3n_{1}+n_{2}$ respectively. Figure 7 shows the spectrum computed numerically for the potential strengths $\lambda=3,4$ respectively for the directions (1,1), (3,1); and size $L=201$ along the z-direction. Both spectra are symmetric with respect to $E=0$ and feature a $E=0$ flatband.
@@
Figure 8: Rescaled participation entropy $\mathcal{S}(k,\lambda)/\ln(L)$ of the $E=0$ flatband eigenstates as a function of $k,\lambda$ for the potential (19). System size $L=201$ along the potential $z$, $\xi=5$. Potential direction $(1,1):z=n_{1}+n_{2}$ (a) and $(3,1):z=3n_{1}+n_{2}$ (b). .
@@
Figure 9: Participation entropy in units of $\ln(L)$ as a function of the potential strength $\lambda$ for the potential (19) with $\xi=5$, for different system sizes $L$ along the z-direction. Potential directions (a): $(1,1),z=n_{1}+n_{2}$ and (b): $(3,1),z=3n_{1}+n_{2}$. Solid lines correspond to momentum $k=0$ while dashed lines correspond to $k=3\pi/4$.
@@
Similarly to the case of Aubry-André potential, the participation entropy (16) for flatband eigenstates for different momenta $k$ reveals a localization-delocalization transition along the z-direction as shown in Fig. 8. For the Aubry-André potential the transition was also revealed by the analytical argument, however no such argument is known for a generic bounded potential, like Eq. (19). Therefore to confirm the transition we analyzed the finite size scaling with the increase of the system size $L$ along the z-direction. The results are shown in Fig. 9 for the two orientations of the potential, two representative momenta $k=0,3\pi/4$, and multiple system sizes $L$ along the z-direction. Participation entropy curves overlap on top of each other or increase to saturation for small values of $\lambda$ with the system size, while the curves decrease with system size for large values of $\lambda$. This strongly supports a localization-delocalization transition as a function of $k$ and $\lambda$ for the arctan potential, and similarly to the AA case. The presence of localized and delocalized eigenstates in the flatband for the AA (17) and arctan (19) potentials suggest that this might be a generic feature of the models (1) with bounded potentials.
@@
Figure 10: Profiles of the flatband eigenstates as a function of $z$ for different momentum and different orientation of the potential (19) for $\lambda=1.5$. The profiles are more localized for larger momenta for other equal parameters. The direction $(3,1)$ leads to larger localization volume than direction $(1,1)$ for the same set of parameters.
@@
Figure 10 shows profiles of the flat band eigenstates for two potential directions $(1,1),(3,1)$ and several momenta. The profiles are more localized for $(1,1)$ direction as compared to $(3,1)$ direction, as well as for larger momenta $k$.
@@ IV.3 Unbounded polynomial potential
The two examples provided above correspond to bounded potentials, i.e. $|V(z)|<\infty$. Here we consider an example of an unbounded polynomial potential:
@@
in Hamiltonians (12) for $(1,1):z=n_{1}+n_{2}$ and (III) for $(3,1):z=3n_{1}+n_{2}$.
Figure 11: Part of the band structure contains the flatband $E=0$. The band structure is obtained by diagonalizing (12) for subfigure (a) and (III) for subfigure (b) for system size $L=201$ along $z$ for potential (20) with potential strength $\lambda=0.2$.
Figure 11 shows the numerically computed spectrum for system size $201$ along the z-direction and potential strength $\lambda=0.2$. The spectrum is symmetric with respect to the origin and there is a $E=0$ flatband. Interestingly we do not observe an all-bands-flat spectrum in this case in contrast to the linear Wannier-Stark case [29], although the potential is also unbounded in this case.
Figure 12: Profiles of the flatband eigenstates as a function of $z$ for different momenta and different orientations of the unbounded potential (20) at fixed $\lambda=0.2$. The direction $(3,1)$ leads to larger localization volume than direction $(1,1)$ for the same set of parameters.
Unlike the bounded potentials, variation of the potential strength $\lambda$ does not change localization properties of the flatband eigenstates: they are always localized. Still localization of the eigenstates depends weakly on the momentum $k$ and is more pronounced for different direction of the potential as shown in Fig. 12.
@@ V Conclusion
We considered square lattice with tight-binding Hamiltonians and anisotropic onsite potentials that vary along one direction only. If such a Hamiltonian is anti-symmetric under simultaneous reflection and time-reversal a flat band emerges at $E=0$ for odd lattice sizes along the potential direction. This flatband construction extends to other Bravais lattices (see Appendix A), and can be thought of as a generalization of the Wannier-Stark flatbands. The existence of the $E=0$ flatband forbids equipotential hopping, but is otherwise robust to the presence of longer-range hoppings under some mild constraints.
@@
We conjectured the absence of CLS for the $E=0$ flatband considered. The localization properties of the flatband eigenstates were analyzed in the quasi-momenta (defined along the equipotential lattice direction) basis which implies extended wave packet along the equipotential lattice direction. For the two bounded potentials we observe a localization transition in the flatbands eigenstates along the direction of the change of the potential with the increase of the potential strength, while no such transition is observed for unbounded potentials. For the quasiperiodic Aubry-André potential this transition is described analytically using a duality transformation. Understanding whether such transformation exists for other potentials and its identification for other bounded potentials is an open problem.
@@
The Ref. 1 introduced an anti-$\mathcal{PT}$ symmetry which ensures an $E=0$ flatband in non-Bravais lattices. The parity operator associated with the anti-$\mathcal{PT}$ symmetry acts uniformly at each unit cell and cannot be defined for tight-binding Hamiltonians on Bravais lattices. The parity operator introduced in this work is different from that used in Ref. 1: it distinguishes odd and even lattice sites along the potential and is defined for Bravais lattices.
@@
For Bravais lattices, tight-binding Hamiltonians in a DC field, i.e. linear potential, and in the absence of equipotential hopping, all the bands in the spectrum are flat [29]. We observed that for other types of unbounded potentials the Hamiltonian does not support all-band-flat structure, in general. Presence of a flatband combined with dispersive bands can have special importance for non-trivial topology and geometry induced exotic phenomena [38, 39].
Our model can be simulated using quantum XX spins on higher dimensional lattices by tuning longitudinal magnetic field locally that would be an equivalent of the anisotropic potential. Our theoretical analysis can be also implemented in the state-of-art cold atomic devices [40, 41].
@@ Acknowledgements.
We are grateful to Sergej Flach for stimulating discussions. The authors acknowledge the financial support from the Institute for Basic Science (IBS) in the Republic of Korea through the project IBS-R024-D1. AM acknowledges the financial support from the National Science Centre, Poland (Grant No. 2021/03/Y/ST2/00186) within the QuantERA II Programme, which has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733.
@@ Appendix A $E=0$ flatbands for generic $d$-dimensional Bravais lattices with anisotropic potentials
We consider here a generalization of the case of the square lattice from Sec. II to the case of a $d$-dimensional Bravais lattice. We use some of the definitions from our previous work on Wannier-Stark flatbands [29]. The lattice sites $\vec{n}=\sum_{j}n_{j}\vec{a}_{j}$ are indexed by a set of integers $\{n_{j}\}$, and unit cell basis vectors $\vec{a}_{j}$ need not be orthonormal. We consider an anisotropic potential that varies along a lattice direction $\vec{\mathcal{E}}$ only and is constant in the perpendicular lattice directions. We assume that the direction $\vec{\mathcal{E}}$ is commensurate as defined in Ref. 29: one can choose $(d-1)$ linearly independent lattice vectors perpendicular to $\vec{\mathcal{E}}$. Under these conditions the potential can be expressed as a function of the lattice coordinate $z=\vec{\mathcal{E}}\cdot\vec{n}/\mathcal{F}$ only, where the constant $\mathcal{F}$ ensures that $z$ is integer and takes all possible values from $\mathbb{Z}$ [29, 1]. The tight-binding Hamiltonian reads
@@
The coordinates $w_{s}$ are not $z$-independent in general, however they can be expressed as linear combinations of coordinate $z$ and $z$-independent $(d-1)$-dimensional vector $\vec{\eta}$. Each of the components $\eta_{i}$ is integer and a linear function of $\{n_{j}\}$—c.f. Eqs. (10) and (13) in the main text for special cases, and Refs. 29, 1 for the general form of $\vec{\eta}$. In this new coordinate $(z,\vec{\eta})$ the Hamiltonian (21) reads
$\mathcal{H}=\sum_{z,\vec{\eta}}$ $\bigg{[}V(z)\outerproduct{z,\vec{\eta}}{z,\vec{\eta}}-\sum_{\vec{l}}t(l_{z},\vec{\epsilon})\outerproduct{z,\vec{\eta}}{z+l_{z},\vec{\eta}+\vec{\epsilon}}\bigg{]}$ (23)
where $l_{z}=\vec{\mathcal{E}}\cdot\vec{l}/\mathcal{F}$, $\vec{\epsilon}$ is the hopping vector along $\vec{\eta}$ for a fixed $z$ and it is a linear function of $\vec{l}$. Importantly, the Hamiltonian is translationally invariant along the $\vec{\eta}$ directions. Therefore it is diagonalizable with respect to lattice momentum $\vec{k}$ conjugate to $\vec{\eta}$.
@@
We define parity operator which reflects a lattice point with respect to a reference lattice point $\vec{n}=\vec{0}$.
@@
We consider real and anti-symmetric potentials:
@@
Therefore the total Hamiltonian as well as the $\vec{k}$-dependent Hamiltonians are anti-symmetric:
@@
[…]
@@
This condition prohibits equipotential hoppings. Additionally it implies that specific directions $\vec{\mathcal{E}}$ and long-range hopping do not support an $E=0$ flatband. For examples on a square lattice: (i) potential change along $\vec{\mathcal{E}}/\mathcal{F}=(1,1)$ implies $z=n_{1}+n_{2}$. If we include the long-range diagonal hoppings $\vec{l}=\pm(1,1)\implies l_{z}=\pm 2$ violates (26); (ii) A tilted potential along $\vec{\mathcal{E}}/\mathcal{F}=(2,1)$ implies $z=2n_{1}+n_{2}$ and the nearest-neighbor hoppings $\vec{l}=(\pm 1,0)$ violates (26).
The anti-symmetry of the Hamiltonian $\mathcal{H}(\vec{k})$ implies that the eigenvalues come in pairs $(E(\vec{k}),-E(\vec{k}))$ with corresponding eigenvectors $(|\psi(\vec{k})\rangle$, $\langle\vec{k}|\mathcal{T}\cdot\mathcal{P}|\vec{k}\rangle|\psi(\vec{k})\rangle)$ for each $\vec{k}$. Together with a center of reflection $z=0$ we have an odd number of degrees of freedom for a fixed momentum, therefore the block Hamiltonian $\mathcal{H}(\vec{k})$ contains odd number of independent eigenvectors. Therefore there is an eigenvalue which has to be its own negation: $E(\vec{k})=-E(\vec{k})$, and therefore there is an eigenvalue $E(\vec{k})=0$ for all $\vec{k}$, i.e. a flatband.
Note that only $\mathcal{P}$ is enough to flip the sign of the potential. But we use extra $\mathcal{T}$. Individually $\mathcal{T}$ or $\mathcal{P}$ flips the sign of momentum $\vec{k}$, but together they do not. This is the reason that the anti-symmetry of the whole Hamiltonian transfers to the anti-symmetry of block Hamiltonians $\mathcal{H}(\vec{k})$ for every momentum $\vec{k}$—which is necessary for the existential proof of the flatband.
Impossibility of the anti-$\mathcal{P}\mathcal{T}$ symmetric flatband in triangular lattice with nearest-neighbor hopping
In the main text Sec. III we provided examples of square lattice Hamiltonians with potential along the main diagonal and direction (3,1). Here we show that for the nearest-neighbor hopping the triangular lattice geometry does not obey the anti-symmetry condition (26). For the triangular lattice, if we align one of the unit cell basis vectors $\vec{a}_{1}$ along a Cartesian axis, the other one will be titled by angle $\pi/3$: $\vec{a}_{2}\cdot\vec{a}_{1}=\cos(\pi/3)$. Therefore in the orthonormal Cartesian basis $(\hat{e}_{1}=\vec{a}_{1},\hat{e}_{2})$ a lattice vector reads
@@
A nearest neighbor Hamiltonian with a potential reads
$|\displaystyle\mathcal{H}=\sum_{n_{1},n_{2}}\bigg{[}V(z)\outerproduct{n_{1},n_{2}}{n_{1},n_{2}}-$ | -
$-\sum_{p=\pm 1}\outerproduct{n_{1},n_{2}}{n_{1}+p,n_{2}}+\outerproduct{n_{1},n_{2}}{n_{1},n_{2}+p}+$
+|$+\outerproduct{n_{1},n_{2}}{n_{1}+p,n_{2}-p}\bigg{]}\;.$ (28)
Here the coordinate $z$ is taken as a linear combination of $n_{1}$ and $n_{2}$
$z=\alpha_{1}n_{1}+\alpha_{2}n_{2}$ (29)
with coprime coefficients $\alpha_{1},\alpha_{2}$—as ensured by the proper choice of $\mathcal{F}$ depending on $\vec{\mathcal{E}}$ [29]. Therefore for the hopping present in Hamiltonian (28)
@@
For the validity of the anti-symmetry condition (26), $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{1}-\alpha_{2}$ have to be odd numbers simultaneously, which is impossible.

### v2

@@ IV.2 Arctan potential
Figure 7: (a), (b) The band structure around the flat band at $E=0$, obtained by diagonalizing (12) and (III.1), respectively. The potential (19) parameters are $\lambda=4$(a) and $\lambda=3$(b), $\xi=5$, and system size $L=201$ along $z$.
We consider the following trigonometric potential in Hamiltonians (12) and (III.1),
$V(z)=\lambda\arctan\left(\frac{z}{\xi}\right),$ (19)
@@
with $\xi=5$, for two different directions $(1,1)$ with $z=n_{1}+n_{2}$ and $(3,1)$ with $z=3n_{1}+n_{2}$ respectively. Figure 7 shows the spectrum computed numerically for the potential strengths $\lambda=4$ and $\lambda=3$ for directions $(1,1)$ and $(3,1)$, respectively; the system size is $L=201$ along the $z$ direction. Both spectra are symmetric with respect to $E=0$ and feature a $E=0$ flat band.
@@
Figure 8: Rescaled participation entropy $\mathcal{S}(k,\lambda)/\ln(L)$ of the $E=0$ flat-band eigenstates as a function of $k,\lambda$ for the potential (19). System size $L=201$ along the potential $z$, $\xi=5$. Potential direction $(1,1):z=n_{1}+n_{2}$ (a) and $(3,1):z=3n_{1}+n_{2}$ (b).
@@
Figure 9: Participation entropy in units of $\ln(L)$ as a function of the potential strength $\lambda$ for the potential (19) with $\xi=5$, for different system sizes $L$ along the $z$ direction. Potential directions (a) $(1,1),z=n_{1}+n_{2}$ and (b) $(3,1),z=3n_{1}+n_{2}$. Solid lines correspond to momentum $k=0$ while dashed lines correspond to $k=3\pi/4$.
@@
Similarly to the case of the Aubry-André potential, the participation entropy (16) for flat-band eigenstates for different momenta $k$ reveals a localization-delocalization transition along the $z$ direction as shown in Fig. 8. For the Aubry-André potential, the transition was also revealed by the analytical argument, however, no such argument is known for a generic bounded potential, like Eq. (19). Therefore, to confirm the transition, we analyzed the finite-size scaling with the increase of the system size $L$ along the $z$ direction. The results are shown in Fig. 9 for the two orientations of the potential, two representative momenta $k=0,3\pi/4$, and multiple system sizes $L$ along the $z$ direction. Participation entropy curves overlap on top of each other or increase to saturation for small values of $\lambda$ with the system size, while the curves decrease with system size for large values of $\lambda$. This strongly supports a localization-delocalization transition as a function of $k$ and $\lambda$ for the arctan potential, similarly to the AA case. The presence of localized and delocalized eigenstates in the flat band for the AA (17) and arctan (19) potentials suggest that this might be a generic feature of the models (1) with bounded potentials.
@@
Figure 10: Profiles of the flat-band eigenstates as a function of $z$ for different momentums and different orientations of the potential (19) for $\lambda=15$. For identical parameters, profiles are more localized for larger momenta $k$. The direction $(3,1)$ leads to larger localization volume than direction $(1,1)$ for the same set of parameters.
@@
Figure 10 shows profiles of the flat band eigenstates for two potential directions $(1,1),(3,1)$ and several momenta. The profiles are more localized for the $(1,1)$ direction as compared to the $(3,1)$ direction, as well as for larger momenta $k$.
@@ IV.3 Unbounded polynomial potential
The two examples provided above correspond to bounded potentials, i.e., $|V(z)|<\infty$. Here we consider an example of an unbounded polynomial potential:
@@
in Hamiltonians (12) for $(1,1):z=n_{1}+n_{2}$ and (III.1) for $(3,1):z=3n_{1}+n_{2}$.
Figure 11: Part of the band structure contains the flat band $E=0$. The band structure is obtained by diagonalizing (12) for (a) and (III.1) for (b) for system size $L=201$ along $z$ for potential (20) with potential strength $\lambda=0.2$.
Figure 11 shows the numerically computed spectrum for system size $201$ along the $z$ direction and potential strength $\lambda=0.2$. The spectrum is symmetric with respect to $E=0$ and there is a $E=0$ flat band. Interestingly, we do not observe an all-bands-flat spectrum in this case in contrast to the linear Wannier-Stark case [28], although the potential is also unbounded in this case.
Figure 12: Profiles of the flat-band eigenstates as a function of $z$ for different momenta and different orientations of the unbounded potential (20) at fixed $\lambda=0.2$. The direction $(3,1)$ leads to larger localization volume than direction $(1,1)$ for the same set of parameters.
Unlike the bounded potentials, variation of the potential strength $\lambda$ does not change localization properties of the flat-band eigenstates: they are always localized. Still localization of the eigenstates depends weakly on the momentum $k$ and is more pronounced for different directions of the potential as shown in Fig. 12. The asymptotic localization properties in 2D (3D) Bravais lattices for arbitrary polynomial potential can be estimated using the approximate analytical treatment reported in Sec. II B of Ref. [37], which suggests a superexponential decay along the $z$ axis in our setting.
@@ V Conclusion
We considered square lattice with tight-binding Hamiltonians and anisotropic on-site potentials that vary along one direction only. If such a Hamiltonian is anti-symmetric under simultaneous reflection and time-reversal a flat band emerges at $E=0$ for odd lattice sizes along the potential direction. This flat-band construction extends to other Bravais lattices (see Appendix A), and can be thought of as a generalization of the Wannier-Stark flat bands. The existence of the $E=0$ flat band forbids equipotential hopping, but is otherwise robust to the presence of longer-range hoppings under some mild constraints.
@@
We conjectured the absence of CLS for the $E=0$ flat band considered. The localization properties of the flat-band eigenstates were analyzed in the quasimomenta (defined along the equipotential lattice direction) basis, which implies an extended wave packet along the equipotential lattice direction. For the two bounded potentials, we observe a localization transition in the flat-band eigenstates along the direction of the change of the potential with the increase of the potential strength, while no such transition is observed for unbounded potentials. For the quasiperiodic Aubry-André potential, this transition is described analytically using a duality transformation. Understanding whether such transformation exists for other potentials and its identification for other bounded potentials is an open problem.
@@
The Ref. [30] introduced an anti-$\mathcal{PT}$ symmetry which ensures an $E=0$ flat band in non-Bravais lattices. The parity operator associated with the anti-$\mathcal{PT}$ symmetry acts uniformly at each unit cell and cannot be defined for tight-binding Hamiltonians on Bravais lattices. The parity operator introduced in this paper is different from that used in Ref. [30]: it distinguishes odd and even lattice sites along the potential and is defined for Bravais lattices.
@@
For Bravais lattices, tight-binding Hamiltonians in a dc field, i.e., linear potential, and in the absence of equipotential hopping, all the bands in the spectrum are flat [28]. We observed that for other types of unbounded potentials, the Hamiltonian does not support an all-bands-flat structure, in general. Presence of a flat band combined with dispersive bands can have special importance for nontrivial topology and geometry induced exotic phenomena [38, 39].
A possible setting to implement Hamiltonians considered in our works could be single magnon excitations in quantum XX spins on higher dimensional lattices. A single magnon Hamiltonian would map onto a single particle tight-binding problem. A longitudinal magnetic field with a specific spatial profile applied to a single magnon in the spin system would then correspond to the anisotropic potential. One experimental platform is cold atomic devices based on Rydberg atoms, e.g., 87Rb. The ground and excited states of the Rydberg atom are equivalent, respectively, to the spin-down (or absence of particle) and spin-up (or presence of particle) states [40, 41, 42]. The detuning parameter provides extra potential energy to the excited Rydberg state compared to its ground state. Anisotropically varying the detuning parameter for the array of 87Rb Rydberg atoms could mimic the desired longitudinal magnetic Z potential for the XX spins [40, 41, 42]. A linear anisotropic potential along the diagonal direction of a square lattice using spatially resolved microwave spectroscopy has been implemented recently [43]. By changing the Helmholtz coil configuration, we might change the linear potential to desired anisotropic bounded or unbounded potentials [43]. A superconducting processor was also used to implement linear gradient potential [44], which may be modified to realize arbitrary potentials.
@@ Acknowledgements.
We are grateful to Sergej Flach for stimulating discussions. The authors acknowledge the financial support from the Institute for Basic Science (IBS) in the Republic of Korea through Project No. IBS-R024-D1. A.M. acknowledges financial support from the National Science Centre, Poland (Grant No. 2021/03/Y/ST2/00186) within the QuantERA II Programme, which has received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 101017733.
@@ Appendix A $E=0$ flat bands for generic $d$-dimensional Bravais lattices with anisotropic potentials
We consider here a generalization of the case of the square lattice from Sec. II to the case of a $d$-dimensional Bravais lattice. We use some of the definitions from our previous work on Wannier-Stark flat bands [28]. The lattice sites $\vec{n}=\sum_{j}n_{j}\vec{a}_{j}$ are indexed by a set of integers $\{n_{j}\}$ and unit cell basis vectors $\vec{a}_{j}$ need not be orthonormal. We consider an anisotropic potential that varies along a lattice direction $\vec{\mathcal{E}}$ only and is constant in the perpendicular lattice directions. We assume that the direction $\vec{\mathcal{E}}$ is commensurate as defined in Ref. [28]: one can choose $(d-1)$ linearly independent lattice vectors perpendicular to $\vec{\mathcal{E}}$. Under these conditions, the potential can be expressed as a function of the lattice coordinate $z=\vec{\mathcal{E}}\cdot\vec{n}/\mathcal{F}$ only, where the constant $\mathcal{F}$ ensures that $z$ is integer and takes all possible values from $\mathbb{Z}$ [28, 30]. The tight-binding Hamiltonian reads
@@
The coordinates $w_{s}$ are not $z$ independent, in general, however, they can be expressed as linear combinations of coordinate $z$ and $z$-independent $(d-1)$-dimensional vector $\vec{\eta}$. Each of the components $\eta_{i}$ is integer and a linear function of $\{n_{j}\}$—cf. Eqs. (10) and (13) in the main text for special cases, and Refs. [28, 30] for the general form of $\vec{\eta}$. In this new coordinate $(z,\vec{\eta})$, the Hamiltonian (21) reads
$\mathcal{H}=\sum_{z,\vec{\eta}}$ $\bigg{[}V(z)\outerproduct{z,\vec{\eta}}{z,\vec{\eta}}-\sum_{\vec{l}}t(l_{z},\vec{\epsilon})\outerproduct{z,\vec{\eta}}{z+l_{z},\vec{\eta}+\vec{\epsilon}}\bigg{]},$ (23)
where $l_{z}=\vec{\mathcal{E}}\cdot\vec{l}/\mathcal{F}$, $\vec{\epsilon}$ is the hopping vector along $\vec{\eta}$ for a fixed $z$ and is a linear function of $\vec{l}$. Importantly, the Hamiltonian is translationally invariant along the $\vec{\eta}$ directions. Therefore, it is diagonalizable with respect to lattice momentum $\vec{k}$ conjugate to $\vec{\eta}$:
@@
We define a parity operator which reflects a lattice point with respect to a reference lattice point $\vec{n}=\vec{0}$:
@@
We consider real and antisymmetric potentials:
@@
Therefore, the total Hamiltonian as well as the $\vec{k}$-dependent Hamiltonians are antisymmetric,
@@
$\matrixelement{\vec{k}}{\mathcal{T}\cdot\mathcal{P}}{\vec{k}}\cdot\mathcal{H}(\vec{k})=-\mathcal{H}(\vec{k})\cdot\matrixelement{\vec{k}}{\mathcal{T}\cdot\mathcal{P}}{\vec{k}},$
@@
This condition prohibits equipotential hoppings. Additionally, it implies that specific directions $\vec{\mathcal{E}}$ and long-range hopping do not support an $E=0$ flat band. For examples on a square lattice: (i) potential change along $\vec{\mathcal{E}}/\mathcal{F}=(1,1)$ implies $z=n_{1}+n_{2}$. If we include the long-range diagonal hoppings, $\vec{l}=\pm(1,1)\implies l_{z}=\pm 2$ violates (26); (ii) a tilted potential along $\vec{\mathcal{E}}/\mathcal{F}=(2,1)$ implies $z=2n_{1}+n_{2}$ and the nearest-neighbor hoppings $\vec{l}=(\pm 1,0)$ violates (26).
The antisymmetry of the Hamiltonian $\mathcal{H}(\vec{k})$ implies that the eigenvalues come in pairs $(E(\vec{k}),-E(\vec{k}))$ with corresponding eigenvectors $(|\psi(\vec{k})\rangle$, $\langle\vec{k}|\mathcal{T}\cdot\mathcal{P}|\vec{k}\rangle|\psi(\vec{k})\rangle)$ for each $\vec{k}$. Together with a center of reflection $z=0$, we have an odd number of degrees of freedom for a fixed momentum, therefore the block Hamiltonian $\mathcal{H}(\vec{k})$ contains an odd number of independent eigenvectors. Therefore, there is an eigenvalue which has to be its own negation: $E(\vec{k})=-E(\vec{k})$, and therefore there is an eigenvalue $E(\vec{k})=0$ for all $\vec{k}$, i.e., a flat band.
Note that only $\mathcal{P}$ is enough to flip the sign of the potential. But we use extra $\mathcal{T}$. Individually, $\mathcal{T}$ or $\mathcal{P}$ flips the sign of momentum $\vec{k}$, but together they do not. This is the reason that the antisymmetry of the whole Hamiltonian transfers to the antisymmetry of block Hamiltonians $\mathcal{H}(\vec{k})$ for every momentum $\vec{k}$—which is necessary for the existential proof of the flatband.
A.1 Impossibility of the anti-$\mathcal{P}\mathcal{T}$ symmetric flat band in triangular lattice with nearest-neighbor hopping
In the main text, Sec. III.1, we provided examples of square lattice Hamiltonians with potential along the main diagonal and direction (3,1). Here we show that for the nearest-neighbor hopping, the triangular lattice geometry does not obey the antisymmetry condition (26). For the triangular lattice, if we align one of the unit cell basis vectors $\vec{a}_{1}$ along a Cartesian axis, the other one will be titled by angle $\pi/3$: $\vec{a}_{2}\cdot\vec{a}_{1}=\cos(\pi/3)$. Therefore, in the orthonormal Cartesian basis $(\hat{e}_{1}=\vec{a}_{1},\hat{e}_{2})$, a lattice vector reads
@@
A nearest-neighbor Hamiltonian with a potential reads
$|\displaystyle\mathcal{H}=\sum_{n_{1},n_{2}}\bigg{[}V(z)\outerproduct{n_{1},n_{2}}{n_{1},n_{2}}-\sum_{p=\pm 1}\outerproduct{n_{1},n_{2}}{n_{1}+p,n_{2}}$ |
+|$+\outerproduct{n_{1},n_{2}}{n_{1},n_{2}+p}+\outerproduct{n_{1},n_{2}}{n_{1}+p,n_{2}-p}\bigg{]}\;.$ (28)
Here the coordinate $z$ is taken as a linear combination of $n_{1}$ and $n_{2}$,
$z=\alpha_{1}n_{1}+\alpha_{2}n_{2},$ (29)
with coprime coefficients $\alpha_{1},\alpha_{2}$—as ensured by the proper choice of $\mathcal{F}$ depending on $\vec{\mathcal{E}}$ [28]. Therefore, for the hopping present in Hamiltonian (28):
@@
For the validity of the antisymmetry condition (26), $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{1}-\alpha_{2}$ have to be odd numbers simultaneously, which is impossible.